# Generating Synthetic Functional Data for Privacy-Preserving GPS Trajectories

### v1

@@ 1 Introduction
With the growing prevalence of personal data disclosure, identity theft, and unauthorized access to personal information, individuals and organizations are increasingly wary of sharing sensitive data. The transformation in data management and perception has enhanced the need for laws to control data usage and to preserve the privacy of individuals behind those data. Governments have invested significant effort in proposing and enacting comprehensive privacy legislation, such as the General Data Protection Regulation (GDPR) introduced by the European Union in 2018 as “the toughest privacy and security law in the world” (Proton Technologies AG, 2024). Moreover, traditional methods of data publishing, which involve releasing possibly identifiable data, are increasingly drawing attention due to the privacy risks that they naturally present. There has been a growing recognition of the need for more sophisticated and robust approaches to data anonymization and publishing. For instance, after showing the vulnerability of data privacy, Sweeney proposed the concept of $k$-anonimity and many techniques arose to ensure data to be $k$-anonymous before publication. Another example is from the work of Dwork et al., who theorized the notion of differential privacy for achieving uninformativeness in data publishing (Dwork and Roth, 2014). Alongside the growing attention to privacy, there remains a strong interest of industries, academia, and research institutes in collecting or acquiring new representative datasets for data-driven policymaking. Undoubtedly, the utility of large amounts of continuously collected data is undeniable, as it allows us to describe phenomena, highlight peculiarities that might otherwise be challenging to uncover, and especially guide decision-makers toward data-driven and informed solutions.
@@ 2.1 State-of-the-art methods for data synthesis
In this context, in the emerging field of computer science, generative Artificial Intelligence (AI) has started providing synthetic data as a natural output of trained models, assessing their utility by including them in the training process, and suggesting that they can be used as a substitute for original data. Among these, Generative Adversarial Networks (GANs) are an example used in various application domains (Goodfellow et al., 2014), as well as Variational Auto-Encoders (VAEs) (Kingma and Welling, 2014). Although the domain of generative AI is predominant, the research on SDG methods has also developed on the parallel track of statistical analysis. Unlike the black-box models typical of NNs, statistical methods inform a model designed for the original data and, by appropriately transforming its input, obtain synthetic data. Examples include random sampling, parametric regression models, and Classification and Regression Trees (CART), as proposed in Nowok et al. (2016), but also Bayesian Network models (Gogoshin et al., 2021), Markov models (Lencastre et al., 2023), and Copula-Based models (DataCebo Inc., 2020). Thus, a primary classification of SDG methods inherits from the research field where their theory is developed. There can be either black-box models of deep-learning approaches, or white-box methods from the statistical field. The method proposed in this research falls into the second group, but in the following there will be reviewed and discussed SDG methods from both categories adhering to this distinction.
@@
As distinct types of data require specifically tailored models, SDG methods can be classified by the type of data to be synthesized, from standard tabular data to more complex structured data like images and time series. This analysis focuses on multivariate GPS trajectory data from smartphones, characterized by groups and sequentiality, i.e., trajectories. Due to this, SDG methods for sequential data are appropriate. Among deep-learning-oriented methods for sequences, it is to be mentioned the DoppelGANger, a GAN architecture proposed by Lin et al. (2020b) and implemented on the open-source platform by Gretel Labs (Lin et al., 2020a; Gretel Labs, 2020). The authors start from a Recurrent Neural Network (RNN), a benchmark in the literature for generating synthetic time series, and modify it to overcome its limitations. The results of this GAN are remarkable due to the improvements over previous architectures, establishing DoppelGANger as one of the benchmarks in the field. It is also worth citing the proposal in the Synthetic Data Vault package by DataCebo (DataCebo Inc., 2020), that is the Conditional Probabilistic Auto-Regressive (CPAR) model (Zhang et al., 2022). By combining the principles of autoregressive models with probabilistic NN, it achieves improved accuracy in capturing complex dependencies within sequential data and generating realistic synthetic samples. It is a joint-modelling procedure, aimed at specifying the joint distribution of the data and sampling new data from the fitted probability distribution. Other deep-learning methods for sequential data synthesis include TimeGAN (Yoon et al., 2019), RCGAN (Esteban et al., 2017), and Sig-WGAN (Liao et al., 2023). In contrast, within the realm of statistical learning, it is important to mention methods that utilize Markov models, proven to outperform AI algorithms in the study by Lencastre et al. (2023).
@@ 3.1 Functional data representation
The initial dataset consists of GPS measurements that record the positions during the day-to-day displacements of anonimized users at specific times. The fundamental GPS signal is composed of (at least) the two coordinates for the spatial location, $c_{1}$ and $c_{2}$, and the absolute time for the temporal indication, $t$, leading to points of the form $P=(c_{1},c_{2},t)$. The signals describing a single trip of a user are grouped in trajectories. A trajectory describes the unique spatiotemporal movement of a single user through a sequence of GPS data points ordered with respect to the absolute time, $T_{i}=\{P_{1i},\dots,P_{N_{i}i}\}\quad\forall i=1,\dots,N$. One can hence see the initial dataset as a set of trajectories $\{T_{i}\}_{i=1,\dots,N}$, each composed of sequences of points $P_{ji}=(c_{1},c_{2},t)_{ji}$ for location and time identification.
@@
$c^{\prime}_{1ji}=\frac{\max c_{1\cdot\cdot}-c_{1ji}}{\max c_{1\cdot\cdot}-\min c_{1\cdot\cdot}}\in[0,1]\quad\forall j=1,\dots,N_{i}\quad\forall i=1,\dots,N$ (1)
Consequently, the original dataset is represented by a series of constants: $\min c_{1}$, $\max c_{1}$, $\min c_{2}$, $\max c_{2}$, $\min t$ and $\max t$, and by the set of trajectories $\{T^{\prime}_{i}\}_{i=1,\dots,N}$ formed by the points $P^{\prime}_{ji}=(c_{1}^{\prime},c_{2}^{\prime},t_{e}^{\prime})_{ji}$.
@@ 3.2 Choice of the appropriate functional metric space
Let $I\subset\mathbb{R}$ be any interval domain. As every interval can be mapped into the unitary one, for the sake of simplicity, let $I=[0,1]$. Consider the space of $p$-dimensional absolutely continuous [^1] HERE I WILL WRITE THE DEF OF ABSOLUTELY CONTINUOUS FUNCTION functions over $I$:
@@ 3.3 Proposed synthetic data generation algorithm
The FDASynthesis method is implemented as a two-step approach, where first distances between curves are computed, and then the proper synthetis phase occurs. An outline of the methodology is presented in Algorithm 1.
@@
/* Iteratively synthetize new data: */
@@
Formally, given a dataset of $N$ functions $\{\bm{f}_{i}\}_{i=1,\dots,N}\subset\mathcal{F}(I)$, the reference function of interest $\bm{f}_{i}$ is chosen and its $K$ closest functions $\{\bm{f}_{ik}\}_{k=1,\dots,K}$ are determined. Note that, in what follows, the index $i=1,\dots,N$ refers to objects associated to the fixed reference function, while the indices $\{ik\}_{k=1,\dots,K}$ represent the $K$ indices of the neighbor functions ordered with respect to their distance from the reference function. Then, let $\bm{d}_{i}=[d_{i1},\dots,d_{iK}]$ represents the $K$-dimensional vector of distances between the reference function $i$ and its $K$ nearest functions, whose elements are denoted as:
$d_{ik}=\textbf{D}[i,i{k}]=\mathcal{D}(\bm{f}_{i},\bm{f}_{ik})\quad\forall k=1,\dots,K$ (12)
@@
In line with the choice of reference space for the functions, the weighted average is computed as the weighted mean of the SRVF of the $K$ neighbor curves and then mapped back to the original space with the correct translation of the starting point, as expressed in Equation 8 and Equation 4. Given the SRVF transformations of the neighbors $\{\bm{q}_{ik}\}_{k=1,\dots,K}$, the average SRVF is computed precisely as:
$\bm{q}_{i}^{s}=\operatorname*{arg\,min}_{\bm{q}\in L^{2}(I)}\sum_{k=1}^{K}p_{k}\min_{\gamma_{k}\in\Gamma}\lVert\bm{q}-(\bm{q}_{ik}\circ\gamma_{k})\sqrt{\dot{\gamma_{k}}}\rVert^{2}$ (15)

### v2

@@ 1 Introduction
With the growing prevalence of personal data disclosure, identity theft, and unauthorized access to personal information, individuals and organizations are increasingly wary of sharing sensitive data. The transformation in data management and perception has enhanced the need for laws to control data usage and to preserve the privacy of individuals behind those data. Governments have invested significant effort in proposing and enacting comprehensive privacy legislation, such as the General Data Protection Regulation (GDPR) introduced by the European Union in 2018 as “the toughest privacy and security law in the world” (Proton Technologies AG, 2024). Moreover, traditional methods of data publishing, which involve releasing possibly identifiable data, are increasingly drawing attention due to the privacy risks that they naturally present. There has been a growing recognition of the need for more sophisticated and robust approaches to data anonymization and publishing. Alongside the growing attention to privacy, there remains a strong interest of industries, academia, and research institutes in collecting or acquiring new representative datasets for data-driven policymaking. Undoubtedly, the utility of large amounts of continuously collected data is undeniable, as it allows us to describe phenomena, highlight peculiarities that might otherwise be challenging to uncover, and especially guide decision-makers toward data-driven and informed solutions.
@@ 2.1 State-of-the-art methods for data synthesis
In this context, in the emerging field of computer science, generative Artificial Intelligence (AI) has started providing synthetic data as a natural output of trained models, assessing their utility by including them in the training process, and suggesting that they can be used as a substitute for original data. Among these, Generative Adversarial Networks (GANs) are an example used in various application domains (Goodfellow et al., 2014), as well as Variational Auto-Encoders (VAEs) (Kingma and Welling, 2014). Although the domain of generative AI is predominant, the research on SDG methods has also developed on the parallel track of statistical analysis. Unlike the black-box models typical of neural networks, statistical methods inform a model designed for the original data and, by appropriately transforming its input, obtain synthetic data. Examples include random sampling, parametric regression models, and Classification and Regression Trees (CART), as proposed by Nowok et al. (2016), but also Bayesian Network models (Gogoshin et al., 2021), Markov models (Lencastre et al., 2023), and Copula-Based models (DataCebo Inc., 2020). Thus, a primary classification of SDG methods inherits from the research field where their theory is developed. There can be either black-box models of deep-learning approaches, or white-box methods from the statistical field. The method proposed in this research falls into the second group, but in the following there will be reviewed and discussed SDG methods from both categories adhering to this distinction.
@@
As distinct types of data require specifically tailored models, SDG methods can be classified by the type of data to be synthesized, from standard tabular data to more complex structured data like images and time series. This analysis focuses on multivariate GPS data from smartphones, characterized by groups and sequentiality, i.e., trajectories. Due to this, SDG methods for sequential data are appropriate. Among deep-learning-oriented methods for sequences, it is to be mentioned the DoppelGANger, a GAN architecture proposed by Lin et al. (2020b) and implemented on the open-source platform by Gretel Labs (Lin et al., 2020a; Gretel Labs, 2020). The authors start from a Recurrent Neural Network (RNN), a benchmark in the literature for generating synthetic time series, and modify it to overcome its limitations. The results of this GAN are remarkable due to the improvements over previous architectures, establishing DoppelGANger as one of the benchmarks in the field. It is also worth citing the proposal in the Synthetic Data Vault package by DataCebo (DataCebo Inc., 2020), that is the Conditional Probabilistic Auto-Regressive (CPAR) model (Zhang et al., 2022). By combining the principles of autoregressive models with probabilistic NN, it achieves improved accuracy in capturing complex dependencies within sequential data and generating realistic synthetic samples. It is a joint-modelling procedure, aimed at specifying the joint distribution of the data and sampling new data from the fitted probability distribution. Other deep-learning methods for sequential data synthesis include TimeGAN (Yoon et al., 2019), RCGAN (Esteban et al., 2017), and Sig-WGAN (Liao et al., 2023). In contrast, within the realm of statistical learning, it is important to mention methods that utilize Markov models, proven to outperform AI algorithms in the study by Lencastre et al. (2023).
@@ 3.1 Functional data representation
The initial dataset consists of GPS measurements that record the positions during the day-to-day displacements of anonimized users at specific times. The fundamental GPS signal is composed of (at least) the two coordinates for the spatial location, $c_{1}$ and $c_{2}$, and the absolute time for the temporal indication, $t$, leading to points of the form $P=(c_{1},c_{2},t)$. The signals describing a single trip of a user are grouped in trajectories. A trajectory describes the unique spatiotemporal movement of a single user through a sequence of GPS data points ordered with respect to the absolute time, $T_{i}=\{P_{i1},\dots,P_{iN_{i}}\}\quad\forall i=1,\dots,N$. One can hence see the initial dataset as a set of trajectories $\{T_{i}\}_{i=1,\dots,N}$, each composed of sequences of points $P_{ij}=(c_{1},c_{2},t)_{ij}$ for location and time identification.
@@
$c^{\prime}_{1ij}=\frac{\max c_{1\cdot\cdot}-c_{1ij}}{\max c_{1\cdot\cdot}-\min c_{1\cdot\cdot}}\in[0,1]\quad\forall j=1,\dots,N_{i}\quad\forall i=1,\dots,N$ (1)
Consequently, the original dataset is represented by a series of constants: $\min c_{1}$, $\max c_{1}$, $\min c_{2}$, $\max c_{2}$, $\min t_{e}$ and $\max t_{e}$, and by the set of trajectories $\{T^{\prime}_{i}\}_{i=1,\dots,N}$ formed by the points $P^{\prime}_{ij}=(c_{1}^{\prime},c_{2}^{\prime},t_{e}^{\prime})_{ij}$.
@@ 3.2 Choice of the appropriate functional metric space
Let $I\subset\mathbb{R}$ be any interval domain. As every interval can be mapped into the unitary one, for the sake of simplicity, let $I=[0,1]$. Consider the space of $p$-dimensional absolutely continuous functions over $I$:
@@ 3.3 Proposed synthetic data generation algorithm
The FDASynthesis method is implemented as a two-step approach, where first distances between curves are computed, and then the proper synthesis phase occurs. An outline of the methodology is presented in Algorithm 1.
@@
/* Iteratively synthesize new data: */
@@
Formally, given a dataset of $N$ functions $\{\bm{f}_{i}\}_{i=1,\dots,N}\subset\mathcal{F}(I)$, the reference function of interest $\bm{f}_{i}$ is chosen and its $K$ closest functions $\{\bm{f}_{k|i}\}_{k=1,\dots,K}$ are determined. Note that, in what follows, the index $i=1,\dots,N$ refers to objects associated to the fixed reference function, while the indices $\{k|i\}_{k=1,\dots,K}$ represent the $K$ indices of the neighbor functions ordered with respect to their distance from the reference function $i$. Then, let $\bm{d}_{i}=[d_{i1},\dots,d_{iK}]$ represents the $K$-dimensional vector of distances between the reference function $i$ and its $K$ nearest functions, whose elements are denoted as:
$d_{ik}=\mathcal{D}(\bm{f}_{i},\bm{f}_{k|i})\quad\forall k=1,\dots,K$ (12)
@@
In line with the choice of reference space for the functions, the weighted average is computed as the weighted mean of the SRVF of the $K$ neighbor curves and then mapped back to the original space with the correct translation of the starting point, as expressed in Equation 8 and Equation 4. Given the SRVF transformations of the neighbors $\{\bm{q}_{k|i}\}_{k=1,\dots,K}$, the average SRVF is computed precisely as:
$\bm{q}_{i}^{s}=\operatorname*{arg\,min}_{\bm{q}\in L^{2}(I)}\sum_{k=1}^{K}p_{k}\min_{\gamma_{k}\in\Gamma}\lVert\bm{q}-(\bm{q}_{k|i}\circ\gamma_{k})\sqrt{\dot{\gamma_{k}}}\rVert^{2}$ (15)